OpenStudy (anonymous):

Find the rate of change of the area of a circle with respect to the radius r. What is the rate when r = 9?

OpenStudy (anonymous):

\[A = \pi r^{2}\] \[\frac{dA}{dr} = \pi \frac{d}{dr}r^{2}\] \[\frac{dA}{dr} = 2\pi r\]